# Continuous Variable Structured Collision Models

### v1

@@ 1 Introduction
Due to their versatility, collision models have become a popular method of investigating a wide variety of phenomena in both Markovian[7, 5, 3, 8] and non-Markovian regimes [9, 10, 11], quantum thermal machines [12, 13, 14, 15, 16, 17, 18, 19, 20] and the behaviour of time crystals [21, 22]. Applications of such models can be found in the field of quantum optics [23, 24, 25, 26], quantum resources and information [27, 28, 29, 30], and many others [31, 32, 33].
A key advantage of these models is the ability to study the thermodynamic properties of open systems [34, 35, 36, 37, 38, 39, 40, 41], allowing for the investigation of the flows of the system’s internal energy, alongside the corresponding thermodynamic heat and work flows. Furthermore, the entropy and entropy production are readily available [42, 43], allowing for the verification of the laws of thermodynamics.
@@
The second example considers spring-like interactions between the system and environment, which may result in divergences when taking the continuous time limit. We find that discrete time propagation must be used in the general case, with the continuous time limit only possible in special cases. We emphasise, however, that these cases are not achievable without the use of a structured environment with at least two oscillators. While this example does not result in a steady state, a method of inducing a steady state through the inclusion of a secondary unstructured environment interacting with the system via a beamsplitter-like interaction is also presented.
@@ 2.1 System and Environment Structure
For simplicity, in this work we consider uncoupled system oscillators of unit mass $m$, with $\hbar=m=1$. Furthermore, the position and momentum operators obey the standard commutation relation $[x_{S_{i}},p_{S_{j}}]=i\delta_{ij}$. Eq. (2) can be expressed in matrix notation as
@@
with $x_{E_{i}}$ denoting the position operator of each environmental oscillator $i$, and $p_{E_{i}}$ the corresponding momentum operator obeying the analogous commutation relations, $[x_{E_{i}},p_{E_{j}}]=i\delta_{ij}$. The Hamiltonians $H_{E}^{I}$ and $H_{SE}$ are then dependent on the choice of internal environmental interactions and system-environment interactions, respectively. The environmental Hamiltonian can also be cast in matrix form, with
@@ 2.2 System Dynamics
with $\langle\dot{R}\rangle_{\text{tot}}=\frac{d}{dt}\langle R_{\text{tot}}\rangle$. The elements of this propagation matrix $D_{\text{tot}}$ can be found directly using the Heisenberg equation for the system and environment position and momentum operators, or equivalently through $D_{\text{tot}}=\Omega_{N_{\text{tot}}}M_{\text{tot}}$, where
@@
The detailed derivation of these superoperators, along with the explicit expressions of $\mathcal{L}^{S}_{2},\mathcal{L}^{SE}_{2}$ and $\mathcal{L}^{S/SE}_{2}$, can be found in A. We have introduced the notation $[A]_{S}$ to denote the partial trace of the environment through the discarding of elements of the covariance matrix outside of those corresponding to the system, retaining only matrix elements $A_{ij}$ for which $1\leq i\leq 2N_{S}$ and $1\leq j\leq 2N_{S}$. Here, $D_{S}$ is a matrix of dimension $2N_{S}$ describing the dynamics of the free Hamiltonian of the system such that $D_{S}=\Omega_{N_{S}}M_{S}$. $D_{SE}$ is then a matrix of dimension $2N_{\text{tot}}$, such that $D_{SE}=\Omega_{N_{\text{tot}}}M_{SE}$.
As we will see throughout the examples in Sec. 3, $D_{SE}$ is dependent on the choice of system-environment interaction Hamiltonian. The terms relating to the free evolution of the environment do not survive the tracing out of the environmental unit; instead the internal environmental couplings only affect the system through the initial state $\sigma_{\text{tot}}(t)$ and the second order $D_{SE}$ terms.
@@ 2.3 Thermodynamic Quantities
The change in the system’s internal energy $\Delta U(t)$, the heat $\Delta Q(t)$ exchanged with the environment and the associated thermodynamic work $\Delta W(t)$ are natural quantities of interest that can be extracted from the dynamics described in Sec. 2.2. These can be expressed through
@@ 3.1 Initial States
Due to the thermal nature of the environmental units and the system’s initial state, each element of $\langle R_{\text{tot}}\rangle$ is identically zero and remains zero throughout the propagation as $\langle\dot{R}_{\text{tot}}\rangle$ is described through Eq. (18), i.e. as a linear combination of the elements of $\langle R_{\text{tot}}\rangle$. The dynamics of the system can therefore be fully described through the dynamics of its corresponding covariance matrix.
@@ 3.2 Example I: Beamsplitter Interactions
The corresponding Lyapunov equation can then be derived by taking
This then leads to a differential equation in the continuous time limit, i.e. $\delta t\to 0$,
@@
$\langle x_{\hat{E}}^{2}\rangle_{\text{sq}}=\frac{1}{2e^{2r}\omega_{E}}\coth\left(\frac{\omega_{E}}{2T_{\hat{E}}}\right)$ (53)
@@
$\langle p_{\hat{E}}^{2}\rangle_{\text{sq}}=\frac{e^{2r}\omega_{E}}{2}\coth\left(\frac{\omega_{E}}{2T_{\hat{E}}}\right),$ (54)
@@
We shall consider two examples in the following sections: one in which the system interacts exclusively with the mode with the smallest frequency, that is, the center of mass $\tilde{\omega}_{\text{CoM}}$, and one in which it interacts with the mode with the largest frequency, $\tilde{\omega}_{\text{Max}}$.
@@ 3.2.1 Center of Mass
Immediately, one can see agreement with the first law of thermodynamics (see Eq. (30)). For the case of $\omega_{S}=\omega_{E}$, one can recognise an inability to produce or absorb work in the center of mass case, explained through the commutator $[H_{S}+H_{E},H_{SE}]$. Following Ref. [34], the condition $[H_{S}+H_{E},H_{SE}]=0$ is sufficient to ensure $\dot{W}(t)=0$. The fact that $\dot{U}(t),\dot{Q}(t)$ and $\dot{W}(t)$ can be expressed as the product of a common function $f(t)$ and the frequencies $\omega_{S}$ and $\omega_{E}$ is reminiscent of similar results for a continuously operated Otto cycle (see for instance Ref. [13]). This is a consequence of the linearity of the coupling between the harmonic oscillators in the environment and the system. For the $\omega_{S}=\omega_{E}$ case, one finds in the steady state
@@
Due to the lack of squeezing or change in the effective temperature, for $\omega_{S}=\omega_{E}$ the dynamics of the system follows the intuitive path: $T_{E}>T_{S}$ leads to the system heating until a steady state at temperature $T_{E}$ is reached; $T_{E}<T_{S}$ leads to the system cooling until a steady state at temperature $T_{E}$ and the $T_{E}=T_{S}$ case produces no changes within the system. As there is no contribution from any other normal mode frequencies, and $\tilde{\omega}_{E_{\rm{CoM}}}$ does not depend on $\lambda$, it is clear in this case that the structure within the environment does not affect the dynamics or steady state of the system.
@@
3.2.2 Largest Normal Mode Frequency
The second case of interest is when the system interacts solely with the mode corresponding with the largest frequency $\tilde{\omega}_{\rm{Max}}$. This corresponds to the normal mode frequency $\tilde{\omega}_{E_{m^{\prime}}}$, where $m^{\prime}$ is chosen to be the value of $m$ for which Eq. 40 is maximised. One finds the maximum normal mode frequency to be
@@ 3.2.2 Largest Normal Mode Frequency
In this regime, the system’s dynamics can be mapped to that of a quantum harmonic oscillator interacting with environmental units prepared in a squeezed thermal state with effective temperature $T_{\hat{E}}$ (see Eq. (58)) and squeezing parameter $r$ (see Eq. (55)). We plot the dependence of the effective temperature $T_{\hat{E}}$ on the environment’s internal coupling strength $\lambda_{I}$ for $N_{E}=3,4,5$ in Fig. 2. The temperature $T_{\hat{E}}$ is always monotonically decreasing with $\lambda_{I}$, such that $T_{\hat{E}}>T_{E}$ for $\lambda_{I}<0$ and $T_{\hat{E}}\leq T_{E}$ for $\lambda_{I}\leq 0$. The dependence of the effective temperature on $N_{E}$ is very small since the maximum frequency $\tilde{\omega}_{\text{Max}}$ depends only weakly on the number of environment oscillators. In the absence of internal interactions within the environment, i.e. $\lambda_{I}\to 0$, one recovers the center of mass dynamics with $T_{\hat{E}}=T_{E}$, analogous to those for beamsplitter-like interactions via a collision model between the system and single oscillator environmental units prepared in a thermal state.
Figure 2: Effective temperature of the environmental units $T_{\hat{E}}$ interacting with the system as a function of the internal coupling strength $\lambda_{I}$ within the environment. The blue (solid), red (dashed) and green(dotted) lines represent the effective temperature of an environmental unit composed of $3$, $4$ and $5$ oscillators respectively, where the system interacts solely with the maximum normal mode frequency of each unit. Parameters: $\omega_{E}=1$, $T_{E}=1$, $\tilde{\gamma}_{E_{m^{\prime}}}=0.5$.
@@
Figure 3: The steady-state energy $H_{\rm{Steady}}$ (left) and work $\dot{W}$ (right) applied to a system interacting with the largest normal mode frequency of the environmental units, each composed of $4$ oscillators prepared in a thermal state at a temperature $T_{E}$ as a function of the internal coupling strength $\lambda_{I}$ (solid line). The horizontal dashed line represents the energy of a thermal state $E_{\rm eq}$. The vertical lines act as a guide to the eye, with the dashed line at $\lambda_{I}=0$ and the dotted line at $\lambda_{I}=\lambda_{I_{C}}$. The steady state internal energy is independent of internal coupling, such that $\dot{U}=0$ for all $\lambda_{I}$, while the heat flow can be found through $\dot{Q}=-\dot{W}$. Parameters: $\omega_{S}=\omega_{E}=1$, $T_{E}=1$, $\tilde{\gamma}_{E_{m^{\prime}}}=0.5$.
@@ 3.3 Example II: Spring-like Interactions
The dynamics of the system can then be described in the form of Eq. (45). We consider two approaches to study these dynamics. The first is by propagation through discrete time steps wherein $\delta t$ is kept finite through
@@
Some of the terms in Eq. (80) may lead to difficulties in taking the continuous limit, as the method of rescaling used in Sec. 3.2 may lead to divergences. In this model, the rescaling method would require taking $\tilde{\lambda}_{E_{m}}^{2}\delta t\to{\rm const.}$ as $\delta t\to 0$. However when doing so, the term $D_{S}^{\text{RN}}$ would diverge through its dependence on $\bar{\omega}_{S}$, due to the divergence of $\tilde{\lambda}_{E_{m}}$.
@@ 3.3.2 Continuous Time
It is possible to choose interaction constants $\lambda_{E_{i}}$ such that divergences do not appear, allowing for the continuous-time limit to be taken. Applying the rescaling $\lambda_{E_{i}}\lambda_{E_{j}}\delta t\to\Lambda_{E_{ij}}$ as $\delta t\to 0$, can in general produce $\mathcal{L}_{1}(\sigma_{\text{tot}})\to\infty$ however this can be avoided provided $\sum_{i=1}^{N_{E}}\lambda_{E_{i}}=0$.
@@
To maximise the effects of the environment, we present the dynamics due to interactions of the system with the largest-frequency normal mode, $\tilde{\omega}_{E_{\text{Max}}}$, as described in Eq. (73). In Fig. 6, we plot the second moments, $\langle x^{2}_{S}\rangle$ and $\langle p_{S}^{2}\rangle$, which oscillate around a linearly increasing average value. The almost linear increase in these quantities is reminiscent of a diffusion model, in which the variances increase linearly with respect to time. Similarly to the discrete case, this is due to a persistent energy injection from the environment.
@@
Figure 6: Dynamics of a system interacting with the largest frequency normal mode in the continuous-time limit. The second moments (left) and thermodynamic quantities (right) are plotted as a function of time $t$. The second moments are described by the elements of the covariance matrix, particularly $\langle x_{S}^{2}\rangle$ (blue, solid), $\langle p_{S}^{2}\rangle$ (red, dashed)) and $\langle x_{S}p_{S}+p_{S}x_{S}\rangle$ (green, dotted). The thermodynamic quantities presented are the flow of internal energy $\dot{U}(t)$ (blue, solid), heat $\dot{Q}(t)$ (red, dashed) and work $\dot{W}(t)$ (green, dotted). Parameters: $\omega_{S}=\omega_{E}=1$, $T_{S}=T_{E}=1$, $\lambda_{I}=0.67$, $\tilde{\Lambda}_{E_{m^{\prime}}}=0.5$.

### v2

@@ 1 Introduction
Due to their versatility, collision models have become a popular method of investigating a wide variety of phenomena in both Markovian[7, 5, 3, 8] and non-Markovian regimes [9, 10, 11, 12], quantum thermal machines [13, 14, 15, 16, 17, 18, 19, 20, 21] and the behaviour of time crystals [22, 23]. Applications of such models can be found in the field of quantum optics [24, 25, 26, 27], quantum resources and information [28, 29, 30, 31], and many others [32, 33, 34, 35, 36].
A key advantage of these models is the ability to study the thermodynamic properties of open systems [37, 38, 39, 40, 41, 42, 43, 44], allowing for the investigation of the flow of the system’s internal energy, alongside the corresponding thermodynamic heat and work flows. Furthermore, the entropy and entropy production are readily available [45, 46], allowing for the verification of the laws of thermodynamics.
@@
The second example considers spring-like interactions between the system and environment, which may result in divergences when taking the continuous time limit. We find that discrete time propagation must be used in the general case, with the continuous time limit only possible in special cases. We emphasise, however, that these cases are not achievable without the use of environmental units composed of a minimum of two oscillators. While this example does not result in a steady state, a method of inducing a steady state through the inclusion of a secondary unstructured environment interacting with the system via a beamsplitter-like interaction is also presented.
@@ 2.1 System and Environment Structure
For simplicity, in this work we consider uncoupled system oscillators of unit mass $m$, with $\hbar=m=1$. Furthermore, the position and momentum operators obey the standard commutation relation $x_{S_{i}},p_{S_{j}}]=i\delta_{ij}$. Eq. (2) can be expressed in matrix notation as
@@
with $x_{E_{i}}$ denoting the position operator of each environmental oscillator $i$, and $p_{E_{i}}$ the corresponding momentum operator obeying the analogous commutation relations, $x_{E_{i}},p_{E_{j}}]=i\delta_{ij}$. The Hamiltonians $H_{E}^{I}$ and $H_{SE}$ are then dependent on the choice of internal environmental interactions and system-environment interactions, respectively. The environmental Hamiltonian can also be cast in matrix form, with
@@ 2.2 System Dynamics
with $\langle\dot{R}\rangle_{\text{tot}}=\frac{d}{dt}\langle R\rangle_{\text{tot}}$. The elements of this propagation matrix $D_{\text{tot}}$ can be found directly using the Heisenberg equation for the system and environment position and momentum operators, or equivalently through $D_{\text{tot}}=\Omega_{N_{\text{tot}}}M_{\text{tot}}$, where
@@
The detailed derivation of these superoperators, along with the explicit expressions of $\mathcal{L}^{S}_{2},\mathcal{L}^{SE}_{2}$ and $\mathcal{L}^{S/SE}_{2}$, can be found in A. We have introduced the notation $A]_{S}$ to denote the partial trace of the environment through the discarding of elements of the covariance matrix outside of those corresponding to the system, retaining only matrix elements $A_{ij}$ for which $1\leq i\leq 2N_{S}$ and $1\leq j\leq 2N_{S}$. Here, $D_{S}$ is a matrix of dimension $2N_{S}$ describing the dynamics of the free Hamiltonian of the system such that $D_{S}=\Omega_{N_{S}}M_{S}$. $D_{SE}$ is then a matrix of dimension $2N_{\text{tot}}$, such that $D_{SE}=\Omega_{N_{\text{tot}}}M_{SE}$.
As we will see throughout the examples in Sec. 3, $D_{SE}$ is dependent on the choice of system-environment interaction Hamiltonian. The terms relating to the free evolution of the environment do not survive the tracing out of the environmental unit; instead the internal environmental couplings only affect the system through the initial state $\sigma_{\text{tot}}(t)$ and the $D_{SE}$ terms.
The discrete propagation of the system can then be described through
The continuous time limit of this propagation is obtained by rescaling the coupling constants $g_{E_{i}}$ within $D_{SE}$ before taking the limit of $\delta t\to 0$, as shown in Sec. 3.2, to form a Lyapunov equation. This is not always possible without divergences, however the following conditions are sufficient to ensure a finite Lyapunov equation:
  1. 1.
    The second order term $\mathcal{L}_{2}$ must be a homogeneous function of the coupling constants $g_{E_{i}}$, where $g_{E_{i}}$ represents the coupling strength between the system and each oscillator $i$ within the environmental unit.
  2. 2.
    The first order corrections to the system’s free evolution appearing in $\mathcal{L}_{1}$ (see Eq. (25)) must be zero: $D_{SE}\sigma_{\text{tot}}(t)]_{S}=[\sigma_{\text{tot}}(t)D_{SE}^{\intercal}]_{S}=0$.
In this work, we consider quadratic coupling Hamiltonians for which condition 1 is always satisfied. In cases for which condition (2) is satisfied, we find that the terms corresponding to the first order expansion in $\delta t$, $\mathcal{L}_{1}(\sigma_{\text{tot}}(t))$, describe the free evolution of the system, while the second order terms, $\mathcal{L}_{2}(\sigma_{\text{tot}}(t))$, result from the interaction of the system with the environment.
@@ 2.3 Thermodynamic Quantities
The change in the system’s internal energy $\Delta U(t)$, the heat $\Delta Q(t)$ exchanged with the environment and the associated thermodynamic work $\Delta W(t)$ are natural quantities of interest [48, 16, 46, 37, 49, 47] that can be extracted from the dynamics described in Sec. 2.2. These can be expressed through
@@ 3.1 Initial States
Due to the thermal nature of the environmental units and the system’s initial state, each element of $\langle R\rangle_{\text{tot}}$ is identically zero and remains zero throughout the propagation as $\langle\dot{R}\rangle_{\text{tot}}$ is described through Eq. (18), i.e. as a linear combination of the elements of $\langle R\rangle_{\text{tot}}$. The dynamics of the system can therefore be fully described through the dynamics of its corresponding covariance matrix.
@@ 3.2 Example I: Beamsplitter Interactions
The corresponding Lyapunov equation can then be derived via Eq. (27), leading to a differential equation in the continuous time limit, i.e. $\delta t\to 0$,
@@
[…]
@@
[…]
@@
We shall consider two examples in the following sections: one in which the system interacts exclusively with the center of mass mode frequency $\tilde{\omega}_{\text{CoM}}=\omega_{E}$, and one in which it interacts with the normal mode frequency corresponding to the largest wavevector within the Brillouin zone and the shortest wavelength supported by the environmental unit.
@@ 3.2.1 Center of Mass
Immediately, one can see agreement with the first law of thermodynamics (see Eq. (31)). For the case of $\omega_{S}=\omega_{E}$, one can recognise an inability to produce or absorb work in the center of mass case, explained through the commutator $H_{S}+H_{E},H_{SE}]$. Following Ref. [37], the condition $H_{S}+H_{E},H_{SE}]=0$ is sufficient to ensure $\dot{W}(t)=0$. The fact that $\dot{U}(t),\dot{Q}(t)$ and $\dot{W}(t)$ can be expressed as the product of a common function $f(t)$ and the frequencies $\omega_{S}$ and $\omega_{E}$ is reminiscent of similar results for a continuously operated Otto cycle (see for instance Ref. [14]). This is a consequence of the linearity of the coupling between the harmonic oscillators in the environment and the system. For the $\omega_{S}=\omega_{E}$ case, one finds in the steady state
@@
Due to the lack of squeezing or change in the effective temperature, for $\omega_{S}=\omega_{E}$ the dynamics of the system follows the intuitive path: $T_{E}>T_{S}$ leads to the system heating until a steady state at temperature $T_{E}$ is reached; $T_{E}<T_{S}$ leads to the system cooling until a steady state at temperature $T_{E}$ and the $T_{E}=T_{S}$ case produces no changes within the system. As there is no contribution from any other normal mode frequencies, and $\tilde{\omega}_{E_{\rm{CoM}}}$ does not depend on $\lambda_{I}$, it is clear in this case that the structure within the environment does not affect the dynamics or steady state of the system.
@@
3.2.2 Largest wavevector within the Brillouin zone
The second case of interest is when the system interacts solely with the normal mode frequency $\tilde{\omega}_{\rm{Max}}$ corresponding to the largest wavevector within the Brillouin zone. This corresponds to the shortest wavelength mode with the frequency $\tilde{\omega}_{E_{m^{\prime}}}$, where $m^{\prime}$ is chosen to be the value of $m$ for which the $\sin^{2}\left[\frac{\pi}{N_{E}}(m-1)\right]$ term in Eq. (41) is maximised. This frequency is found to be
@@ 3.2.2 Largest wavevector within the Brillouin zone
In this regime, the system’s dynamics can be mapped to that of a quantum harmonic oscillator interacting with environmental units prepared in a squeezed thermal state with effective temperature $T_{\hat{E}}$ (see Eq. (58)) and squeezing parameter $r$ (see Eq. (55)). This is chosen as a case of particular interest as it allows for control over the effective temperature of the environment. We plot the dependence of the effective temperature $T_{\hat{E}}$ on the environment’s internal coupling strength $\lambda_{I}$ for $N_{E}=3,4,5$ in Fig. 2. The temperature $T_{\hat{E}}$ is always monotonically decreasing with $\lambda_{I}$, such that $T_{\hat{E}}>T_{E}$ for $\lambda_{I}<0$ and $T_{\hat{E}}\leq T_{E}$ for $\lambda_{I}\leq 0$. The dependence of the effective temperature on $N_{E}$ is very small as $\tilde{\omega}_{\text{Max}}$ depends only weakly on the number of environment oscillators. In the absence of internal interactions within the environment, i.e. $\lambda_{I}\to 0$, one recovers the center of mass dynamics with $T_{\hat{E}}=T_{E}$, analogous to those for beamsplitter-like interactions via a collision model between the system and single oscillator environmental units prepared in a thermal state.
Figure 2: Effective temperature of the environmental units $T_{\hat{E}}$ interacting with the system as a function of the internal coupling strength $\lambda_{I}$ within the environment. The blue (solid), red (dashed) and green(dotted) lines represent the effective temperature of an environmental unit composed of $3$, 44 and 55 oscillators respectively, where the system interacts solely with the normal mode frequency of the environmental units $\tilde{\omega}_{E_{\text{Max}}}$ corresponding to the largest wavevector within the Brillouin zone. Parameters: $\omega_{E}=1$, $T_{E}=1$, $\tilde{\gamma}_{E_{m^{\prime}}}=0.5$.
@@
Figure 3: The steady-state energy $H_{\rm{Steady}}$ (left) and work $\dot{W}$ (right) applied to a system interacting with the normal mode frequency of the environmental units $\tilde{\omega}_{E_{\text{Max}}}$ corresponding to the largest wavevector within the Brillouin zone. Each unit is composed of $4$ oscillators prepared in a thermal state at a temperature $T_{E}$ as a function of the internal coupling strength $\lambda_{I}$ (solid line). The horizontal dashed line represents the energy of a thermal state $E_{\rm eq}$. The vertical lines act as a guide to the eye, with the dashed line at $\lambda_{I}=0$ and the dotted line at $\lambda_{I}=\lambda_{I_{C}}$. The steady state internal energy is independent of internal coupling, such that $\dot{U}=0$ for all $\lambda_{I}$, while the heat flow can be found through $\dot{Q}=-\dot{W}$. Parameters: $\omega_{S}=\omega_{E}=1$, $T_{E}=1$, $\tilde{\gamma}_{E_{m^{\prime}}}=0.5$.
@@
This scheme lends itself well to physical implementation, such as through trapped ion platforms using Coulomb crystals [52]. One can see from Eq. (73) that this model is identical for all even $N_{E}$, with the exception of the special case $N_{E}=2$. Therefore, one only requires $N_{E}=4$ to achieve the maximum deviation from $\omega_{E}$ for a given $\lambda_{I}$. Experimentally, this can be achieved by ‘resetting’ an environmental unit of $4$ oscillators to the initial state between interactions to avoid preparing a large number of units, provided the reset time $\tau_{\text{Reset}}$ is faster than the relaxation time of the system $\tau_{\text{Relax}}$, i.e. $\tau_{\text{Reset}}\ll\tau_{\text{Relax}}$. We emphasise, however, that precise engineering of the environmental units is not crucial to mimic a squeezed bath and that a high-frequency mode coupled to the system’s degrees of freedom would be sufficient.
@@ 3.3 Example II: Spring-like Interactions
The dynamics of the system can then be described in the form of Eq. (27). We consider two approaches to study these dynamics. The first is by propagation through discrete time steps wherein $\delta t$ is kept finite. One can begin with
@@
Some of the terms in Eq. (80) may lead to difficulties in taking the continuous limit, as the method of rescaling used in Sec. 3.2 may lead to divergences. In this model, the rescaling method would require taking $\tilde{\lambda}_{E_{m}}^{2}\delta t\to{\rm const.}$ as $\delta t\to 0$. However when doing so, the term $D_{S}^{\text{RN}}$ would diverge through its dependence on $\bar{\omega}_{S}$, due to the divergence of $\tilde{\lambda}_{E_{m}}$. We remark, however, that using a finite $\delta t$, as in the discrete evolution, would not lead to any mathematical issue. As the continuous time limit is not possible in the general case, we instead propagate the system through Eq. (16) to avoid the approximation due to the Taylor expansion in $\delta t$.
@@ 3.3.2 Continuous Time
It is possible to choose interaction constants $\lambda_{E_{i}}$ such that divergences do not appear, allowing for the continuous-time limit to be taken [53]. Applying the rescaling $\lambda_{E_{i}}\lambda_{E_{j}}\delta t\to\Lambda_{E_{ij}}$ as $\delta t\to 0$, can in general produce $\mathcal{L}_{1}(\sigma_{\text{tot}})\to\infty$ however this can be avoided provided $\sum_{i=1}^{N_{E}}\lambda_{E_{i}}=0$.
@@
To maximise the effects of the environment, we present the dynamics due to interactions of the system with the normal mode frequency $\tilde{\omega}_{E_{\text{Max}}}$ described in Eq. (73). In Fig. 6, we plot the second moments, $\langle x^{2}_{S}\rangle$ and $\langle p_{S}^{2}\rangle$, which oscillate around a linearly increasing average value. The almost linear increase in these quantities is reminiscent of a diffusion model, in which the variances increase linearly with respect to time. Similarly to the discrete case, this is due to a persistent energy injection from the environment.
@@
Figure 6: Dynamics of a system interacting with the normal mode frequency $\tilde{\omega}_{E_{\text{Max}}}$ corresponding to the largest wavevector within the Brillouin zone in the continuous-time limit. The second moments (left) and thermodynamic quantities (right) are plotted as a function of time $t$. The second moments are described by the elements of the covariance matrix, particularly $\langle x_{S}^{2}\rangle$ (blue, solid), $\langle p_{S}^{2}\rangle$ (red, dashed)) and $\langle x_{S}p_{S}+p_{S}x_{S}\rangle$ (green, dotted). The thermodynamic quantities presented are the flow of internal energy $\dot{U}(t)$ (blue, solid), heat $\dot{Q}(t)$ (red, dashed) and work $\dot{W}(t)$ (green, dotted). Parameters: $\omega_{S}=\omega_{E}=1$, $T_{S}=T_{E}=1$, $\lambda_{I}=0.67$, $\tilde{\Lambda}_{E_{m^{\prime}}}=0.5$.